# KiDS-Legacy: Redshift distributions and their calibration

### v1

@@ 6.1.4 $\left[{\textbf{D}}\right]:$ Impact of stellar contamination
We investigate the significance of the contribution of stars to the redshift distributions from SKiLLS by running our calibration pipeline assuming perfect stellar rejection, and comparing the resulting distributions/biases to our fiducial run. These results are presented in Table 5. We find that the difference in bias with and without the stellar contamination is of similar order as the uncertainty in the bias estimates between realisations: the maximal difference in the recovered bias is $|\Delta\langle\delta_{z}\rangle|=0.0078$ in the sixth tomographic bin, which is a roughly $2\sigma$ deviation from the fiducial bias uncertainty (which itself is a lot smaller than the conservative final uncertainty, see Sect. 6.1.8).
6.1.5 $\left[{\textbf{E}}\right]:$ Calibration-side weighting
In previous KiDS analyses, redshift calibration has always been performed without weights utilised on the calibration side of the SOM/direct calibration process. This is primarily because the spectroscopic calibration fields were (and largely remain to be) not wide-field shear fields. However, as discussed in Sect. 4, in KiDS-Legacy we implement additional weighting on the calibration side in an effort to mitigate systematic biases: shape-measurement weights, and prior-volume weights. We test the impact of removing these weights one at a time in Sects. 6.1.6 and 6.1.7. For closer compatibility with previous work, however, here we also test the redshift calibration process when not including either of these weights. The results are presented in Table 5.
In terms of bias, we find that the inclusion of the weighting (i.e. Run ID [A]) produces smaller biases in some bins than the unweighted scenario (i.e. Run ID [E]). These shifts are partly larger than the uncertainty estimates. Bins two, four, and five see reductions in the absolute value of the bias (going from [E] to [A]) at the $2-3\sigma$ level. However, the opposite effect is true in some other bins: bins one and three see a $2\sigma$ increase in absolute value of the bias when including the calibration side weights. Of note is the mechanism for bias to change on the redshift distributions without corresponding change in the estimated redshift distribution means. The weighting on the calibration side influences the effective redshift distribution of the wide-field sample (i.e. the truth) through the gold-weight, which itself is modified by the relative up- and down-weighting of calibration sources by the shape and prior volume weights.
@@ 6.1.5 $\left[{\textbf{E}}\right]:$ Calibration-side weighting
Overall, the weights on the calibration side can be seen to have a non-negligible impact on the redshift distribution biases. The individual impact of the shape-measurement weights and prior-volume weights are discussed in the following sections. Given the changing nature of the wide-field sample $N(z)$ under the calibration-side weighting, we therefore opt not to include these changes in bias as a systematic component in our cosmological analysis. Rather we instead reserve both sets of redshift distributions and weighted source samples for separate cosmological analyses. While the redshift distributions are not directly comparable and the analysis presented here does not strongly favour one approach, the cosmological parameters should still be highly consistent for both scenarios. In practice we use the maximally realistic setup for our fiducial case: including both prior volume weights and shape weights on the calibration side.
6.1.6 $\left[{\textbf{F}}\right]:$ Impact of shape-measurements
We next test the influence on the redshift distributions biases when implementing weighting on the calibration side using only shape-measurement weights. We compare these shape-weight-only biases to those from both the fiducial and no-calibration-weighting results presented in Table 5.
@@ 6.1.6 $\left[{\textbf{F}}\right]:$ Impact of shape-measurements
After SOM weighting, however, the situation is clearer. After SOM weighting, the redshift distribution biases are significantly improved in all bins except for the first, where the calibration side shape-measurement weights are not able to correct the over-correction of the SOM weighting (discussed in Sect. 6.1.5). Overall, the addition of calibration side shape weights produces a reduction in bias compared to the unweighted result, particularly in the bins most sensitive to cosmic shear.
6.1.7 $\left[{\textbf{G}}\right]:$ Impact of prior-weights
@@
After SOM weighting, we again find biases that are improved in some bins and degraded in others: in the higher redshift tomographic bins we see consistent further reduction in biases after the SOM weighting, however we again see over-correction in the lower redshift bins. Relative to the results when using shape-only weighting we see that the prior-volume-only weighting produces slightly poorer bias recovery, but that the bias is reduced relative to the implementation without any calibration-side weighting in the majority of bins.
[…]
We verify the computation of biases using two different simulated data sets, as a means of estimating the robustness of our calibration procedure to the assumptions inherent to a single simulation. To do this, we apply our $N(z)$ estimation algorithm to samples constructed in the SKiLLS (Sect. 3.1) and MICE2 (Sect. 3.2) data sets, where the only a priori modification to the simulations is to ensure that both cover the same underlying redshift baseline: $1.420.07<z_{\rm true}<$ 1.42. This ensures that this test probes the difference that is attributable to the construction of the simulation for a consistent population of source galaxies, rather than probing differences in bias generated by different samples with different redshift extent. Additionally, we opt to compute the difference between the recovered biases in the regime where we do not include shape-measurements on the calibration side of the computation, as the shape measurement weights are systematically different between the simulations: MICE2 shape measurement weights are synthetic and determined by matching colours to the observed data, rather than being measured from images as in SKiLLS.
Resulting redshift distribution biases for our SKiLLS and MICE2 data sets are presented in Table 5. The results show that, for our analysis without calibration side weights and only including the prior volume weights, we find consistent redshift distribution bias parameters for the two simulations at the level of $|\Delta\delta_{z}|\la 0.01$. This indicates that there is an inherent uncertainty floor in the accuracy to which we can estimate the redshift distribution bias parameters from our simulations, driven by the realism of the simulations themselves. Such an error floor is in reality somewhat conservative, as the MICE2 simulations here are known to lack realism in many regards (not the least of which is the lack of imaging and the use of purely analytic photometric noise realisations). As such, it is expected that these simulations ought to diverge in their realism, with SKiLLS being the more trustworthy reference. Nonetheless, we opt to utilise this $|\Delta\delta_{z}|\la 0.01$ systematic difference in our computation of the redshift distribution bias priors, by implementing this as an uncertainty floor in the prior specification.
@@ 6.2 CC redshift distributions
Similar to the SOM analysis, we first test our calibration methodology on MICE2. Since we need to measure two-point statistics on the simulation, we require a slightly different version from those reported in the section above; the version used here applies the matching algorithm neither on the calibration sample nor the photometric data to ensure that the clustering properties of the resulting gold sample are preserved. In addition, we find that, when comparing the shift-fit values $D_{\rm z}$ (Eq. 5) to the SOM bias (Eq. 3), it is preferential to compute the SOM bias as the difference in the median of the redshift distribution instead of the mean, i.e. defining
@@
Table 6: Different redshift bias estimates per tomographic bin obtained from the ensemble of MICE2 CC realisations. The values listed here are the median SOM bias, the mean and standard deviation of the shift-fit parameter obtained by fitting the CCs with the true redshift distribution, followed by fitting with the $N_{\rm SOM}(z)$, and finally the difference between the SOM bias and the corresponding shift-fit parameter.
@@
2 $\phantom{-}0.058\pm 0.010$ $\phantom{-}0.060\pm 0.006$ $-0.002\pm 0.012$
3 $\phantom{-}0.049\pm 0.010$ $\phantom{-}0.056\pm 0.006$ $-0.007\pm 0.011$
@@
5 $-0.004\pm 0.010$ $-0.016\pm 0.019$ $\phantom{-}0.012\pm 0.022$
6 $-0.022\pm 0.012$ $-0.085\pm 0.071$ $\phantom{-}0.063\pm 0.072$
 [^14] The results in this table are not directly comparable to Table 5 as the source samples are inherently different. $\langle D_{\rm z}^{\rm SOM}\rangle$ is duplicated from Table 7 for comparison.
@@ 6.2.1 CC measurements
As described in Sect. 5.4, we do not use the clustering redshift distributions that we measure directly from MICE2. Instead, we first apply the noise adaptation scheme and add the level of intrinsic scatter that we find in the data by fitting the $f$-term (Eq. 6) on the data and create 100 realisations of the intrinsic scatter. The mean and scatter of these realisations are shown in Fig. 10. They closely trace the underlying true redshift distribution of the simulated KiDS-Legacy data (green line).
Figure 10: Comparison of the ensemble of MICE2 realisations and their best-fit solutions when using the SOM redshift distributions as model for the shift-fit. The black data points indicate the mean and standard deviation (encapsulating the added intrinsic scatter) of the CC measurement realisations, the green and blue lines represent the true and SOM redshift distributions, respectively. The blue line and shaded area (median and 68 % confidence interval) is the $N_{\rm SOM}(z)$ after applying the shift-fit parameter value in each realisation.
@@ 6.2.2 CCs fitted with true redshift distributions
First, we need to verify that our updated measurement process and fitting procedure are able to accurately reproduce the true redshifts of the MICE2 galaxies. Therefore, we fit the CC measurements with the true redshift distribution to check whether the resulting shift parameter $D_{\rm z}^{\rm true}$ is consistent with zero in all bins. We compared a number of different correlation measurement scales and found that excluding the smallest scales ($r<0.5\leavevmode\nobreak\ \mathrm{Mpc}$) results in the least biased shift parameters, whereas setting the upper limit to $1.5\leavevmode\nobreak\ \mathrm{Mpc}$ still maintains a good level of signal-to-noise in the correlation amplitude. Therefore, we choose $0.5<r\leq 1.5\leavevmode\nobreak\ \mathrm{Mpc}$ as fiducial measurement scale for KiDS-Legacy.
The ensemble of parameter values from all realisations of this setup, fitted with the true redshift distribution, is listed in Table 7 and shown by the green data points in Fig. 11. Most importantly, all shift-fit values are consistent with zero and $|\langle D_{\rm z}^{\rm true}\rangle|\,<0.01$ (except for bin six), demonstrating that our new methodology is able to produce unbiased estimates of the underlying true redshift distribution. In the first four tomographic bins, where most of the redshift distributions are covered by the wide area samples 2dFLenS, BOSS, and GAMA, the scatter in $D_{\rm z}^{\rm true}$ is $0.005$. In bins five and six, which are dominated by the CCs from DESI, the scatter is significantly larger ($0.017$ and $0.044$) with the largest shift of $\langle D_{\rm z}^{\rm true}\rangle=-0.015$ in bin six.
We find that our fitting approach with the additional error term, accounting for the intrinsic scatter, is able to (on average) reproduce the input value from the measurements of the data. The scatter of the $f$-values between realisations is similar to the uncertainty of the input value. The distribution of goodness-of-fit values is consistent with $\chi^{2}/{\rm dof}=1$ in all bins, as we expect given the ability of the model to increase the magnitude of the uncertainty.
Table 7: Summary of the parameters and goodness of fit obtained from the shift-fits on the data and the 100 MICE2 realisations with noise adaptation. The uncertainties quoted for MICE2 refer to the standard deviation of the realisations, not the error of the mean.
@@
Data SOM $n(z)$ $D_{\rm z}$ $\phantom{-}0.000\pm 0.007$ $\phantom{-}0.028\pm 0.006$ $\phantom{-}0.021\pm 0.006$ $\phantom{-}0.035\pm 0.010$ $-0.018\pm 0.028$ $-0.145\pm 0.091$
@@
[…]
MICE2 realisations True $n(z)$ $\langle D_{\rm z}\rangle$ $\phantom{-}0.003\pm 0.005$ $\phantom{-}0.002\pm 0.005$ $\phantom{-}0.005\pm 0.005$ $-0.004\pm 0.007$ $-0.008\pm 0.017$ $-0.015\pm 0.044$
@@
SOM $n(z)$ $\langle D_{\rm z}\rangle$ $\phantom{-}0.014\pm 0.007$ $\phantom{-}0.060\pm 0.006$ $\phantom{-}0.056\pm 0.006$ $\phantom{-}0.015\pm 0.006$ $-0.016\pm 0.019$ $-0.085\pm 0.071$
@@
$\langle\chi^{2}_{\rm dof}\rangle$ $\phantom{-0}0.97\pm 0.06$ $\phantom{-0}0.98\pm 0.08$ $\phantom{-0}0.97\pm 0.07$ $01\pm 0.1$ 0.10 $\phantom{-0}0.97\pm 0.10$ $\phantom{-0}1.05\pm 0.15$
151515 $D_{z}$: $N(z)$ shift parameter (see Eq. 5) $f$: additive error term (see Eq. 6)
@@ 6.2.3 CCs fitted with SOM redshift distributions
The second step of verifying our pipeline is fitting the mock CCs with the SOM redshift distributions and comparing the shift parameters to the bias in the median SOM redshift ($\delta z_{\rm med}$). Ideally, both values should be in agreement if the difference between the CCs and $N_{\rm SOM}(z)$ can, to first order, be rectified by a simple shift in redshift.
The goodness of fit for these fits is, similar to the case of fitting with the true redshift distributions above, consistent with $\chi^{2}_{\rm dof}=1$. The fitted $f$-values are up to $50\,\%$ larger compared to the previous case but, with the exception of bin one, fully consistent with the input values from the data when considering the uncertainty and the scatter between the realisations (see Table 7). The shift-parameter values are much larger when using the $N_{\rm SOM}(z)$ as fit model and only consistent with zero in bin five; see the red data points in Fig. 11. In bins one to four, $\langle D_{\rm z}^{\rm SOM}\rangle$ is largely positive, indicating that the SOM overestimates the median redshift by about $0.015$ in bins one and four and up to $0.060$ in bin two. In general, the scatter in $D_{\rm z}^{\rm SOM}$ is similar to the ones obtained from the fits with the true redshift distribution.
While these values indicate a large bias in the SOM redshifts, they are perfectly consistent with the SOM bias reported for this specific version of MICE2 (compare Table 6 and the blue confidence regions in Fig. 11). When we take the difference $\delta z_{\rm med}-\langle D_{\rm z}^{\rm SOM}\rangle$ of both bias estimates, the value is consistent with zero considering the scatter of $D_{\rm z}^{\rm SOM}$ between the noise realisations, as indicated by the black data points in Fig. 11. The scatter is about $0.012$ in the first four bins and $0.022$ and $0.072$ in bins five and six, respectively. The amplitude of the difference is less than $0.01$ in all but the last two tomographic bins, of which especially bin six is poorly constrained.
Finally, we can confirm visually in Fig. 10 that applying the shift to the $N_{\rm SOM}(z)$ on MICE2 results in a much better match with the CCs than the unshifted redshifts. Especially in bins two and three, where $\langle D_{\rm z}^{\rm SOM}\rangle$ is large, the peak of the distribution closely matches the realisations of the CCs after shifting. There are some remaining residual differences near the tails of the distributions, which probably explain the increase in the best-fit $f$-values.

### v2

@@ 6.1.5 $\left[{\textbf{E}}\right]:$ Impact of stellar contamination
We investigated the significance of the contribution of stars to the redshift distributions from SKiLLS by running our calibration pipeline assuming perfect stellar rejection, and comparing the resulting $N(z)$ distributions and biases to those from our fiducial run. These results are presented in Table 5. We find that the difference in bias with and without the stellar contamination is of similar order as the uncertainty in the bias estimates between realisations: the maximal difference in the recovered bias is $|\Delta\langle\delta_{z}\rangle|=0.0078$ in the sixth tomographic bin, which is a roughly $2\sigma$ deviation from the fiducial bias uncertainty (which itself is a lot smaller than the conservative final uncertainty; see Sect. 6.1.9).
6.1.6 $\left[{\textbf{F}}\right]:$ Calibration-side weighting
In previous KiDS analyses, redshift calibration has always been performed without weights utilised on the calibration side of the SOM/direct calibration process. This is primarily because the spectroscopic calibration fields were (and largely remain to be) not wide-field shear fields. However, as discussed in Sect. 4, in KiDS-Legacy we implement additional weighting on the calibration side in an effort to mitigate systematic biases: shape-measurement weights, and prior-volume weights. We tested the impact of removing these weights one at a time in Sects. 6.1.7 and 6.1.8. For closer compatibility with previous work, however, we also tested the redshift calibration process when not including either of these weights. The results are presented in Table 5.
In terms of bias, we found that the inclusion of the weighting (i.e. Run ID [A]) produces smaller biases in some bins than the unweighted scenario (i.e. Run ID [E]). These shifts are partly larger than the uncertainty estimates. Bins two, four, and five see reductions in the absolute value of the bias (going from [E] to [A]) at the $2-3\sigma$ level. However, the opposite effect is true in some other bins: bins one and three see a $2\sigma$ increase in absolute value of the bias when including the calibration side weights. Of note is the mechanism for bias to change on the redshift distributions without corresponding change in the estimated redshift distribution means. The weighting on the calibration side influences the effective redshift distribution of the wide-field sample (i.e. the truth) through the gold-weight, which itself is modified by the relative up- and down-weighting of calibration sources by the shape and prior volume weights.
@@ 6.1.6 $\left[{\textbf{F}}\right]:$ Calibration-side weighting
Overall, the weights on the calibration side were found to have a non-negligible impact on the redshift distribution biases. The individual impact of the shape-measurement weights and prior-volume weights are discussed in the following sections. Given the changing nature of the wide-field sample $N(z)$ under the calibration-side weighting, we opted not to include these changes in bias as a systematic component in our cosmological analysis. Rather we instead reserved the two sets of redshift distributions and weighted source samples for separate cosmological analyses. While the redshift distributions are not directly comparable and the analysis presented here does not strongly favour one approach, the cosmological parameters should still be highly consistent for both scenarios. In practice, we use the maximally realistic setup for our fiducial case: including both prior volume weights and shape weights on the calibration side.
6.1.7 $\left[{\textbf{G}}\right]:$ Impact of shape measurements
We next tested the influence on the redshift distributions biases when implementing weighting on the calibration side using only shape-measurement weights. We compared these shape-weight-only biases to those from both the fiducial and no-calibration-weighting results presented in Table 5.
@@ 6.1.7 $\left[{\textbf{G}}\right]:$ Impact of shape measurements
After SOM weighting, however, the situation is clearer: the redshift distribution biases are significantly improved in all bins except for the first, where the calibration side shape-measurement weights are not able to correct the over-correction of the SOM weighting (discussed in Sect. 6.1.6). Overall, the addition of calibration side shape weights produces a reduction in bias compared to the unweighted result, particularly in the bins most sensitive to cosmic shear.
6.1.8 $\left[{\textbf{H}}\right]:$ Impact of prior weights
@@
After SOM weighting, we again found biases that were improved in some bins and degraded in others: in the higher redshift tomographic bins we see consistent further reduction in biases after the SOM weighting; however, we again see over-correction in the lower redshift bins. Relative to the results when using shape-only weighting we see that the prior-volume-only weighting produces slightly poorer bias recovery, but that the bias is reduced relative to the implementation without any calibration-side weighting in the majority of bins.
6.1.9 $\left[{\textbf{I-J}}\right]:$ SKiLLS versus MICE2
We verified the computation of biases using two different simulated datasets, as a means of estimating the robustness of our calibration procedure to the assumptions inherent to a single simulation. To do this, we applied our $N(z)$ estimation algorithm to samples constructed in the SKiLLS (Sect. 3.1) and MICE2 (Sect. 3.2) datasets, where the only a priori modification to the simulations is to ensure that both cover the same underlying redshift baseline: $0.07<z_{\rm true}<1.42$. This ensures that this test probes the difference that is attributable to the construction of the simulation for a consistent population of source galaxies, rather than probing differences in bias generated by different samples with different redshift extent. Additionally, we opted to compute the difference between the recovered biases in the regime where we do not include shape measurements on the calibration side of the computation, as the shape-measurement weights are systematically different between the simulations: MICE2 shape measurement weights are synthetic and determined by matching colours to the observed data, rather than being measured from images as in SKiLLS.
Resulting redshift distribution biases for our SKiLLS and MICE2 datasets are presented in Table 5. For our analysis without calibration side weights and only including the prior volume weights we found consistent redshift distribution bias parameters for the two simulations at the level of $|\Delta\delta_{z}|\la 0.01$. This indicates that there is an inherent uncertainty floor in the accuracy to which we can estimate the redshift distribution bias parameters from our simulations, driven by the realism of the simulations themselves. Such an error floor is in reality somewhat conservative, as the MICE2 simulations here are known to lack realism in many regards (not the least of which is the lack of imaging and the use of purely analytic photometric noise realisations). Thus, it is expected that these simulations ought to diverge in their realism, with SKiLLS being the more trustworthy reference. Nonetheless, we opted to utilise this $|\Delta\delta_{z}|\la 0.01$ systematic difference in our computation of the redshift distribution bias priors, by implementing this as an uncertainty floor in the prior specification.
@@ 6.2 CC redshift distributions
Similar to the SOM analysis, we first tested our calibration methodology on MICE2. Since we needed to measure two-point statistics on the simulation, we required a slightly different version from those reported in the section above; the version used here applies the matching algorithm neither on the calibration sample, nor the photometric data to ensure that the clustering properties of the resulting gold sample are preserved. In addition, we found that, when comparing the shift-fit values $D_{\rm z}$ (Eq. 5) to the SOM bias (Eq. 3), it was preferential to compute the SOM bias as the difference in the median of the redshift distribution instead of the mean, by defining
@@
Table 6: Different redshift bias estimates per tomographic bin obtained from the ensemble of MICE2 $C$ realisations.  [^14] The values listed here are the median SOM bias, the mean and standard deviation of the shift-fit parameter obtained by fitting the CCs with the true redshift distribution, followed by fitting with the $N_{\rm SOM}(z)$, and finally the difference between the SOM bias and the corresponding shift-fit parameter. The results in this table are not directly comparable to Table 5 as the source samples are inherently different. $\langle D_{\rm z}^{\rm SOM}\rangle$ is duplicated from Table 7 for comparison.
@@
2 $\phantom{-}0.058\pm 0.010$ $\phantom{-}0.060\pm 0.006$ $0.002\pm 0.012$
3 $\phantom{-}0.049\pm 0.010$ $\phantom{-}0.056\pm 0.006$ $0.007\pm 0.011$
@@
5 $004\pm 0.010 0.0$$16\pm 0.019 0.0$$2\phantom{-}0.012\pm$ 0.022
[…]
@@ 6.2.1 CC measurements
As described in Sect. 5.4, we did not use the clustering redshift distributions that we measured directly from MICE2. Instead, we first applied the noise adaptation scheme and added the level of intrinsic scatter that we found in the data by fitting the $f$-term (Eq. 6) on the data and created 100 realisations of the intrinsic scatter. The mean and scatter of these realisations are shown in Fig. 10. They closely trace the underlying true redshift distribution of the simulated KiDS-Legacy data (green line).
Figure 10: Comparison of the ensemble of MICE2 realisations and their best-fit solutions when using the SOM redshift distributions as model for the shift-fit. The black data points indicate the mean and standard deviation (encapsulating the added intrinsic scatter) of the CC measurement realisations, the green and blue lines represent the true and SOM redshift distributions, respectively. The red line and shaded area (median and 68 % confidence interval) is the $N_{\rm SOM}(z)$ after applying the shift-fit parameter value in each realisation.
@@ 6.2.2 CCs fitted with true redshift distributions
First, we needed to verify that our updated measurement process and fitting procedure were able to accurately reproduce the true redshifts of the MICE2 galaxies. Therefore, we fit the CC measurements with the true redshift distribution to check whether the resulting shift parameter $D_{\rm z}^{\rm true}$ was consistent with zero in all bins. We compared a number of different correlation measurement scales and found that excluding the smallest scales ($r<0.5\penalty 10000\ \mathrm{Mpc}$) resulted in the least biased shift parameters, whereas setting the upper limit to $1.5\penalty 10000\ \mathrm{Mpc}$ still maintained a good S/N in the correlation amplitude. Therefore, we chose $0.5<r\leq 1.5\penalty 10000\ \mathrm{Mpc}$ as fiducial measurement scale for KiDS-Legacy.
The ensemble of parameter values from all realisations of this setup, fitted with the true redshift distribution, are listed in Table 7 and shown by the green data points in Fig. 11. Most importantly, all shift-fit values are consistent with zero and $|\langle D_{\rm z}^{\rm true}\rangle|\,<0.01$ (except for bin six), demonstrating that our new methodology is able to produce unbiased estimates of the underlying true redshift distribution. In the first four tomographic bins, where most of the redshift distributions are covered by the wide area spectroscopic samples of 2dFLenS, BOSS, and GAMA, the scatter in $D_{\rm z}^{\rm true}$ is $0.005$. In bins five and six, which are dominated by the CCs from DESI, the scatter is significantly larger ($0.017$ and $0.044$) with the largest shift of $\langle D_{\rm z}^{\rm true}\rangle=-0.015$ in bin six.
We found that our fitting approach, with the additional error term that accounted for intrinsic scatter, was able (on average) to reproduce the input value from the measurements of the data. The scatter of the $f$-values between realisations was similar to the uncertainty of the input value. The distribution of goodness-of-fit values was consistent with $\chi^{2}/{\rm dof}=1$ in all bins, as was expected given the model’s ability to increase the magnitude of the uncertainty.
Table 7: Parameters and goodness of fit obtained from $C$ analyses.  [^15] Results were computed from shift-fits on the data and the 100 MICE2 realisations with noise adaptation. The uncertainties quoted for MICE2 refer to the standard deviation of the realisations, not the error of the mean. $D_{z}$: $N(z)$ shift parameter (see Eq. 5) $f$: additive error term (see Eq. 6)
@@
Data SOM $n(z)$ $D_{\rm z}$ $\phantom{-}0.000\pm 0.007$ $\phantom{-}0.028\pm 0.006$ $\phantom{-}0.021\pm 0.006$ $\phantom{-}0.035\pm 0.010$ $0.018\pm 0.028$ $0.145\pm 0.091$
@@
$\chi^{2}_{\rm dof}$ $\phantom{-0}1.44$ $\phantom{-0}1.04$ $\phantom{-0}1.11$ $\phantom{-0}1.03$ $\phantom{-0}1.04$ $\phantom{-0}1.14$
MICE2 realisations True $n(z)$ $\langle D_{\rm z}\rangle$ $\phantom{-}0.003\pm 0.005$ $\phantom{-}0.002\pm 0.005$ $\phantom{-}0.005\pm 0.005$ $004\pm 0.007 0.0$$08\pm 0.017 0.0$$015\pm 0.044$
@@
SOM $n(z)$ $\langle D_{\rm z}\rangle$ $\phantom{-}0.014\pm 0.007$ $\phantom{-}0.060\pm 0.006$ $\phantom{-}0.056\pm 0.006$ $\phantom{-}0.015\pm 0.006$ $016\pm 0.019 0.0$$085\pm 0.071$
@@
$\langle\chi^{2}_{\rm dof}\rangle$ $\phantom{-0}0.97\pm 0.06$ $\phantom{-0}0.98\pm 0.08$ $\phantom{-0}0.97\pm 0.07$ $\phantom{-0}1.01\pm 0.10$ $\phantom{-0}0.97\pm 0.10$ $\phantom{-0}1.05\pm 0.15$
@@ 6.2.3 CCs fitted with SOM redshift distributions
The second step of verifying our pipeline was fitting the mock CCs with the SOM redshift distributions and comparing the shift parameters to the bias in the median SOM redshift ($\delta z_{\rm med}$). Ideally, both values should be in agreement if the difference between the CCs and $N_{\rm SOM}(z)$ can, to first order, be rectified by a simple shift in redshift.
The goodness of fit for these fits is, similar to the case of fitting with the true redshift distributions above, consistent with $\chi^{2}_{\rm dof}=1$. The fitted $f$-values are up to $50\,\%$ larger compared to the previous case but are otherwise, with the exception of bin one, fully consistent with the input values from the data when considering the uncertainty and the scatter between the realisations (see Table 7). The shift-parameter values are much larger when using the $N_{\rm SOM}(z)$ as fit model and only consistent with zero in bin five; see the red data points in Fig. 11. In bins one to four, $\langle D_{\rm z}^{\rm SOM}\rangle$ is largely positive, indicating that the SOM overestimates the median redshift by about $0.015$ in bins one and four and up to $0.060$ in bin two. In general, the scatter in $D_{\rm z}^{\rm SOM}$ is similar to the ones obtained from the fits with the true redshift distribution.
While these values indicated a large bias in the SOM redshifts, they are perfectly consistent with the SOM bias reported for this specific version of MICE2 (compare Table 6 and the blue confidence regions in Fig. 11). When we take the difference $\delta z_{\rm med}-\langle D_{\rm z}^{\rm SOM}\rangle$ of the two bias estimates, the value is consistent with zero considering the scatter of $D_{\rm z}^{\rm SOM}$ between the noise realisations, as indicated by the black data points in Fig. 11. The scatter is about $0.012$ in the first four bins and $0.022$ and $0.072$ in bins five and six, respectively. The amplitude of the difference is less than $0.01$ in all but the last two tomographic bins, of which bin six is particularly poorly constrained.
Finally, we confirmed visually in Fig. 10 that applying the shift to the $N_{\rm SOM}(z)$ on MICE2 results in a much better match with the CCs than the unshifted redshifts. Especially in bins two and three, where $\langle D_{\rm z}^{\rm SOM}\rangle$ is large, the peak of the distribution closely matches the realisations of the CCs after shifting. There are some remaining residual differences near the tails of the distributions, which probably explain the increase in the best-fit $f$-values.